what is 6x(3+2)=18?
Answer by wuwei96815(245)

6x(3+2)=18
First, remove the parenthesis. You get:
18x + 12x = 18
Next, combine like terms. You get:
30x = 18
Divide both sides by 30 in order to isolate the variable x. You get:
x = 18/30 = 3/5 = 0.60